OpenStudy (anonymous):

Solve the equation on the interval [0, 2π). cos x + 2 cos x sin x = 0

OpenStudy (anonymous):

@AnswerMyQuestions

OpenStudy (anonymous):

1. 2 cos x − sin² x = cos² x 2 cos x = sin² x + cos² x 2 cos x = 1 cos x = 1/2

OpenStudy (precal):

I would factor out cosx from both terms cosx(1+2sinx)=0 then set each term to zero cosx=0 1+2sinx=0

OpenStudy (anonymous):

so @victorhernandez what would the answer be?

OpenStudy (precal):

solve for x in each case and then recall that your domain is only from 0 to 2pi

OpenStudy (precal):

I don't know what victorhernandez did but it doesn't look correct unless he rewrote it using identities.

OpenStudy (anonymous):

:3

OpenStudy (anonymous):

That is what i was thinking @precal he didnt write it right

OpenStudy (precal):

http://www.thattutorguy.com/unit-circle-pdf/ you can look up the values on this pdf but any trig table or unit circle will work

OpenStudy (precal):

cosx=0 look at the circle and look at the x values x values are cosine and y values are sine look for (0, number) then look at the degrees or radians where x is 0

OpenStudy (precal):

1+2sinx=0 2sinx=-1 sinx = -1/2 look at the y values and look for -1/2 and that will give you the solutions for sinx=-1/2

OpenStudy (anonymous):

is it pi/2, 7pi/6, 3pi,2,11pi/6

OpenStudy (precal):

let me double check

OpenStudy (precal):

yes that is correct

OpenStudy (anonymous):

YAY! lol Thank you

OpenStudy (precal):

ok but more important did you understand it because then you can do other problems as well.

OpenStudy (anonymous):

yes

OpenStudy (precal):

yw